# A generalization of the Choi isomorphism with application to open quantum systems

### v1

@@ VI.3.6 $t^{2}$-terms
This completes our arguments showing that the GKS-matrix $g(t)$ corresponding to the time evolution of an open quantum system will be positively semi-definite if we consider the expansion $g(t)=g_{0}+g_{1}\,t+g_{2}\,t^{2}+O(t^{3})$ up to second order in $t$. Recall that a non-vanishing $2$nd order contribution $g_{2}\,t^{2}$ to the GKS-matrix corresponds to the case where the semi-group property of the time evolution is violated. Moreover, the restriction to a pure initial state $\rho_{E}(0)$ of the environment is harmless in this respect, since the GKS-matrix $g(t)$ depends linearly on $\rho_{E}(0)$. If $g_{0}+g_{1}\,t+g_{2}\,t^{2}\geq 0$ for pure $\rho_{E}(0)$ then this also holds for all general mixed statistical operators $\rho_{E}(0)$.
@@ VII Summary
Even 70 years after its definition, it seems that the concept of a “completely positive transformation" is still not fully understood mathematically and physically. The same applies to the criterion for complete positivity, which is provided by the Choi isomorphism. This isomorphism depends on the choice of an orthonormal basis $\left(|i\rangle\right)_{0\leq i<N}$ in Hilbert space. An obvious motive in the search for generalizations is therefore the extension of bases of the form $\left(|i\rangle\langle j|\right)_{0\leq i,j<N}$ to arbitrary orthonormal bases $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ in the space of complex $N\times N$-matrices. Such an extension had already existed for some time in the form of the dePillis-Jamiołski isomorphism, but the criterion for complete positivity had to be sacrificed for it. In a recent paper FC24 by Frembs and Cavalcanti, however, this problem has been solved.
@@
implicitly defining the GKS matrix $\left(g^{\alpha\beta}\right)$. For the special case $F_{\alpha}=|j\rangle\langle i|$, the GKS matrix is reduced to the Choi matrix and thus inherits the criterion “${\mathcal{E}}$ is completely positive iff $g\geq 0$.

### v2

@@ VI.3.6 $t^{2}$-terms
This completes our arguments showing that the GKS-matrix $g(t)$ corresponding to the time evolution of an open quantum system will be positively semi-definite if we consider the expansion $g(t)=g_{0}+g_{1}\,t+g_{2}\,t^{2}+O(t^{3})$ up to second order in $t$. Recall that a non-vanishing 22nd order contribution $g_{2}\,t^{2}$ to the GKS-matrix corresponds to the case where the semi-group property of the time evolution is violated. Moreover, the restriction to a pure initial state $\rho_{E}(0)$ of the environment is harmless in this respect, since the GKS-matrix $g(t)$ depends linearly on $\rho_{E}(0)$. If $g_{0}+g_{1}\,t+g_{2}\,t^{2}\geq 0$ for pure $\rho_{E}(0)$ then this also holds for all general mixed statistical operators $\rho_{E}(0)$.
@@ VII Summary
Even 70 years after its definition, it seems that the concept of a “completely positive transformation" is still not fully understood mathematically and physically. The same applies to the criterion for complete positivity, which is provided by the Choi isomorphism. This isomorphism depends on the choice of an orthonormal basis $\left(|i\rangle\right)_{0\leq i<N}$ in Hilbert space. An obvious motive in the search for generalizations is therefore the extension of bases of the form $\left(|i\rangle\langle j|\right)_{0\leq i,j<N}$ to arbitrary orthonormal bases $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ in the space of complex $N\times N$-matrices. Such an extension had already existed for some time in the form of the dePillis-Jamiołski isomorphism, but the criterion for complete positivity had to be sacrificed for it. In the articles PS13 and FC24 , however, this problem has been solved.
@@
implicitly defining the GKS matrix $\left(g^{\alpha\beta}\right)$. For the special case $F_{\alpha}=|j\rangle\langle i|$, the GKS matrix is reduced to the Choi matrix and thus inherits the criterion $\mathcal{E}$} is completely positive iff $g\geq 0$ ".
@@
Acknowledgment
I would like to thank Kyung Hoon Han for valuable comments on earlier versions of this paper, and Frederik vom Ende for a literature reference. I am also grateful to Vittorio Gorini for first-hand information on the early history of the GKSL equation.